OpenStudy (anonymous):

what are the zeros of f(x)=2x^4-2x^2-40?

OpenStudy (mathstudent55):

Let the function equial zero, then solve for x. Try factoring first.

Directrix (directrix):

Zeros occur where f(x) = 0. f(x)=2x^4-2x^2-40 Let f(x) = 0 0 = 2x^4-2x^2-40 0 = 2*( x ^4 - x^2 - 20) @charjay Factor ( x ^4 - x^2 - 20) and post what you get, okay?